# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v1

@@ 3.1 Algorithm Based on Truncation Error Analysis
The basic logic of the diffusion ODE is as follows: When the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as detrimental as the non-noisy steps. These insights have been proved in a previous paper[15] via an experiment that discretizes the inference process by choosing a uniform variant gap in each discretized time step to obtain arrays: $\{\sigma_{i}=\frac{i}{N-1}\sigma_{max};i<64\}$ and $\{t_{i};t(\sigma_{i})\}$. Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$. The experimental results show that the truncation error increases monotonically as the start variant decreases.
[…]
Based on this observation, one researcher [15] proposes a following discrete method:
Eq. 15 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part. Moreover, according to the observation from the paper[17], if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget. Considering the U-Net has 860 million parameters, and the variant of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical. Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
Based on that observation, one study [17] proposes another discrete method:
When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part is larger than that in Karras’s original method.
Furthermore, we should consider the meaning of the score function by rewriting the Eq. 6 as follows:
@@
which means the score function is pointing to the average of all possible images that can cause this noisy image. Naturally, one can surmise that $\nabla_{x}\log q(x_{t})\approx\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage. In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
Based on the aforementioned idea and the experiment results in Section 4, we propose the following discrete method that aims to correct the output in the less noisy stage:
where $\sigma_{stop}$ is defined via the following formula:
$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{N+stop+1}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (19)
which is not zero, as the decode $\beta$-VAE trained with a small KL-term constraint can handle a small amount of noise, and this setting can reduce the one-step inference budget. The visual different between our discrete method and the method proposed by the Karras’ paper is in Fig. 2
Additionally, consider the Eq. 17; the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, which can be achieved by modifying the skip connection and the backbone feature of the U-Net. Previous study[34] adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability. Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the special part of the noisy backbone feature map that lower than a particular frequency. The decorator only applies to the first two skip connections, and the scaler affected the backbone feature is $b_{1}$ and $b_{2}$, while the scaler affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
Also, consider that the normal ResNet[9] can be viewed as an ODE[5] by using the following formula:
@@
The U-Net, as a black box ODE solver, can be regarded as another ODE. Thus, modifying the skip connection and scale backbone feature will further increase its output and the moving distance of the diffusion reverse ODE, which let the start step of using U-Net decorator should be parameterized in 1024 x 1024 image synthesis, as it contributes to the increase of moving distance in the noisy part. This behavior causes the global truncation error in solving ODE, which may let the variables be out of the distribution (OOD). We also visualize the decoded feature in the trajectory of different samplers in Fig. 3.
@@
Figure 3: The comparison between our sampler and the original DPM++1s sampler. The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler. The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
@@
We then visualize the decode trajectory of using different discrete methods by applying the same sampler DPM++ 1s sampler in Fig. 5. The normal time schedule using Free-U will generate a trajectory similar to the method using Karras’ time schedule without using Free-U, which exhibits the evidence of our previous claim.
@@
Figure 4: The comparison between different discrete method when using the same DPM++ 1s Sampler in a few-step inference. The first line exhibit the trajectory of the normal Karras discrete method. The second one exhibit the trajectory of the Karras discrete method + Free-U. The third exhibite the normal discrete method with normal U-Net, and the final one exhibit the normal discrete method with Free-U. And the final line is similiar to the first line.
Our algorithm can be described by the Algorithm 1
@@
Data: number of inference step $N$, initiated noise $\epsilon$, stop step in karras schedule $t_{stop}$, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$
@@
2 $ts=\{t_{i}=(t_{max}^{\frac{1}{p_{1}}}+\frac{i}{N}(t(\sigma_{stop})^{\frac{1}{p_{1}}}-t_{max}^{\frac{1}{p_{1}}}))^{p_{1}};0\leq i<N\}$ $x=\epsilon$ while $t$ in tsts do
3 if $t$ equal to $t_{aug}$ then
4    $freeU(F_{\theta})$;
5
@@ 4    $freeU(F_{\theta})$;
7 $x=F_{\theta}(x)$
@@ 3.2 Information Theory Analysis of the Different Algorithms
A common point of confusion in deep learning at an early age is how neural networks with so many parameters still have generalization ability[1]. The normal statistical model uses an extremely limited number of parameters to ensure that it can concisely describe the phenomenon and capture the essence of the observed data. For this problem, the study theory[8] claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions. Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{n}$. According to the information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can gain access to is $2^{H(D)}$[42]. Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
@@
Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$. The information bottleneck theory[39] shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ project to the $Z$. Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses. As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) [7], which face mode collapse[6] that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint [14]. One study[17] shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information. The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE)[2], in which, when $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bounce of the mutual information constraint will be abated, and there will be an increase in the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$. From this perspective, the diffusion model can be deemed as a hierarchy VAE[25], while progressive distillation[31] increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$. Where $X_{\theta}$ is a noisy feature output by solving ODE. Hence, the PSNR also increases. A recent example of diffusion distillation is SDXL-lightning[19], which adds GAN loss to the progressive distillation training, and its FID remains similar to normal progressive distillation.
@@
The diffusion ODE can be regarded as a Markov process $X_{\theta}(T)-X_{\theta}(T-1)-...-X_{\theta}(t)-...X_{\theta}(0)$, in which $X_{\theta}(t)$ follows the input of the diffusion model predicted by the U-Net, and we have the following Eq. 23:
$I(X_{\theta}(0),X_{\theta}(t))=H(X_{\theta}(0))-H(X_{\theta}(0)|X_{\theta}(t))$ (22)
$\leq H(X(0))-E_{X_{\theta}(t)}[p(X_{\theta}(0)|X_{\theta}(t),$
$*\log p_{\theta}(X_{\theta}(0)|X_{\theta}(t)))],$ (23)
in which $H(X(0))$ is the real $X(0)$ sample from the dataset.
In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$ while enhancing the log likelihood of the final synthesis. Moreover, Eq. 23 shows that if we modify the $X_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase. Most distillation methods that do not increase the PSNR of the output will still face this problem. We evaluate the Precision and Recall of Distribution (PRD)[30] of flash diffusion and our sampler to further validate this claim.

### v2

@@ 3.1 Algorithm Based on Truncation Error Analysis
To design a training-free, few-step inference algorithm, we should first consider the meaning of the original diffusion models’ output, the score function. And then compare it with the distillation model’s output.
By rewriting the eq.˜8 as follows:
@@
We find out that the score function is pointing to the average of all possible images that can cause this noisy image. In an extreme few-step inference case, like one or two NFE inference, the final output will be blurred, since in that step, the score function only points to the centroid of images. Meanwhile, the distillation algorithm, like rectify flow proposed by Liu (2022), will tend to modify the output and let it point to a special sample.
Although the task of modifying the score function to allow it to point to a special sample without additional training is hard, one can presume that the score function $\nabla_{x}\log q(x_{t})$ can be approximated by the conditional score function $\nabla_{x}\log q(x_{t}|x_{0})$ in the least noisy stage, since the possible images are very limited. In this situation, the inference task can be approximated by a pure denoising task, and we do not need further inference if we can modify the U-Net to perform the denoising task without additional training.
Since the changing $\nabla_{x}\log p_{\theta}(x_{t})$ to the $\nabla_{x}\log p_{\theta}(x_{t}|x_{0})$ means removing the blurred part of the result caused by the average, one of the natural method is to enhance the denoise capability of U-Net, which can be achieved by modifying the skip connection and the backbone feature of the U-Net.
Previous study (Si et al. (2024)) adds a scalar, fast Fourier transformation (FFT) and inverse fast Fourier transformation (iFFT) modules to the skip connection, which the backbone feature being modified by the scalar can use to enhance denoising capability. Moreover, the skip connection being modified by the FFT and iFFT can prevent further smoothing and add details by amplifying the low-frequency part of the noisy backbone feature map. The decorator only applies to the first two skip connections. The scaler that affects the backbone feature is $b_{1}$ and $b_{2}$, while the scaler that affects the FFT and iFFT is $s_{1}$ and $s_{2}$.
However, we don’t know which step could be changed simply from the score function to the conditional score function, and one of the vanilla ways is to apply free-U at all time steps. This method has been seen as helpful to slightly modify the score function and make it more similar to the distillation’s output. The previous research from Ma et al. (2024) also shows that compared to the original diffusion model, the diffusion distillation model has small skip connection scalers, since the skip connection itself will abate the result in the block below that layer, while the few-step inference tends to excute a huge transportation from a standard Gaussian distribution to a complex probability distribution, and hence require the skip connection reduce its’ effect that abate the block below it. However, via the mathematical analysis of the diffusion ODE and the experimental result. This vanilla method in the extreme few-step inference of high-resolution images, such as 6 NFE, would not work properly.
Explicitly, consider that the normal ResNet (He et al. (2016)) can be viewed as an ODE (Chen et al. (2018)) by using the following formula:
@@
The U-Net, as a black box ODE solver, can be regarded as another ODE. Thus, modifying the skip connection and scale backbone feature will further increase the norm of the output, increasing the moving distance of the diffusion reverse ODE. In the situation of the few-step inference, it will be a serious problems, since the weakness of the poor discretization method will be amplified by this behaviour. The related mathematic details will be mentioned below. We also visualize the decoded features along the trajectories of different samplers in fig.˜2.
@@
Figure 2: The comparison between our custom DPM++1s sampler and the original DPM++1s sampler. The first line is the result trajectory of the original DPM++1s sampler with Karras’ schedule, the second one is Karras’ schedule + Free-U, and the last one is our sampler. The results show that if we apply Free-U in a few-step sampling without involving additional tricks, the generation result will be degraded.
@@
To further show the influence of the Free-U on the black box ODE solver, we then visualize the decoded trajectory using different discrete methods by applying the same sampler DPM++ 1s sampler in fig.˜3. The standard time schedule using Free-U generates a trajectory similar to that of Karras et al. (2022) time schedule without Free-U, providing evidence that Free-U increases the moving distance at the beginning of the inference, which determines the image layout composition.
@@
Figure 3: The comparison between different discrete methods when using the same DPM++ 1s Sampler in a few-step inference. The first line exhibits the trajectory of the normal Karras discrete method. The second one exhibits the trajectory of the Karras discrete method + Free-U. The third exhibits the normal discrete method with a normal U-Net, and the final one exhibits the normal discrete method with Free-U. And the final line is similar to the first line.
Another method is to parameterize the start step of using the U-Net decorator in an existing solver. However, experimental results show that this Vallina method doesn’t work either. We visualize the 6-step inference result combining the free-U and UniPC solver in fig.˜4.
Figure 4: The result that starts using free-U in the different steps of the UniPC solver.
To address black-box ODE-solving problems, we should identify the characteristics of the diffusion ODE.
The basic logic of the ODE solver is to discretize a continuous curve trajectory into a series of lines. And to achieve the best performance, the ODE solver should determine the length of each discrete line for different ODEs, such as the Fehlberg method (see e.g. Atkinson et al. (2009)) that adjusts the discrete length based on the error per unit stepsize.
For the diffusion ODE, when the moving distance in step $t$ to $t+1$ is large during inference, the weakness of the poor discretion in the noisy stage will not be as fatal as the non-noisy steps. These insights have been proved by Karras et al. (2022) via an experiment that discretizes the inference process with a uniform gap between each timestep. And it obtains an array of steps $\{t_{i}\}$. Then, the researchers calculate the root mean square error (RMSE) between the one-step inference from $t_{i}$ to $t_{i+1}$ and the 200-step inference from $t_{i}$ to $t_{i+1}$. The experimental results show that the truncation error increases monotonically as the noise variance add in the input feature at the time step $t_{i}$ decreases.
[…]
Based on this observation, Karras et al. (2022) proposes a following discrete method:
eq.˜21 controls the variant $\sigma_{i}$ via $p$, and if the $p$ increases, the $\sigma_{i}-\sigma_{i-1}$ will increase in the noisier part and decrease in the less noisy part. Moreover, according to the observation from the Li & Zhang (2025), if we choose $p=7$ and $N=20$, the fourth-to-last step corresponds to the first $\frac{6}{1000}$ perturbation steps, and it will cost $\frac{1}{5}$ inference budget. Considering the U-Net has 860 million parameters, and the variance of the $p(x_{t}|x_{0})$ is approximately $\frac{6}{1000}$, performing four-step ODE inference at this point is illogical. Moreover, the $\beta$-VAE used for decoding is also trained with a small KL-term constraint, which enables it to handle the noise in the latent variable.
Based on that observation, Li & Zhang (2025) proposes another discrete method:
When choosing $p=1.2$, the $h_{i}$ in both the noisy part and the less noisy part are larger than those in Karras’s original method. In such a situation, $t_{i}-t_{i-1}$ will increase in the noisier part. The eq.˜4 shows that $t_{i}$ is proportional to the variance of noise added in $x_{t}$, which can be deemed as the speed of movement at timestep $t_{i}$. The increase of $t_{i}-t_{i-1}$ in the noisy part aligns with the rise in speed, while the rise of $\sigma_{i}-\sigma_{i-1}$ aligns with the increment of the moving distance, hence it can cope with the problem of the original Karras schedule mentioned above.
However, both of those discrete methods have been commonly used in all kinds of images. Our research shows that, even with the same time schedule and UniPC solver, perceptual quality in the few-step inference remains substantially different across image resolution and image type. Hence, the design space of hyperparameters for different ODE solvers has not yet been totally exploited. We find a discrete method, which changes the eq.˜23, a formula with a special design reason mentioned later, by adding one more step to get clean denoised images. And use it to generate images at different resolutions and of different types in a few-step inference. We find that the result in fig.˜5 has a noticeable gap in perceptual quality.
Figure 5: The comparison between the different resolution images and different type images. The first line contains 512 x 512 images generated by Stable Diffusion 1.5, the second line contains 1024 x 1024 images generated by Dreamshaper XL. The last lines contain high-resolution ACGN images generated by Illustrij Evo.
Based on the experimental results, designing a customized training-free algorithm for different image types requires exploring the design space of the ODE solver for the latent diffusion model.
Firstly, the decoder of the latent diffusion model, a $\beta$-VAE trained with a small KL-term constraint, can handle a small amount of noise; thereby, if we choose a time schedule that does not fully denoise the image, we can save further inference budget. Although the latent diffusion model theoretically does not require using the KL term to train an autoencoder, this optimization trick is very prevalent in the training of the diffusion model. The latest representation autoencoders (RAE) also add noise during the training to augment the decoder Zheng et al. (2025). The experimental results show that the decoder’s capacity to handle noisy features varies with image resolution. In the fig.˜6, the second line uses the same time schedule as the first line, which doesn’t denoise totally, but those images in the second line have some blurred parts that harm the details of the images, while the first line doesn’t confront this problem. And the third line output feature with less noise, and has a better results.
Figure 6: The comparison between the different resolution images generated by the same time schedule and ODE solver utilizing the same $\beta$-VAE.
Based on the aforementioned idea and the experiment results in Section 4, we first involve the following discrete method that aims to correct the output in the less noisy stage:
where $\sigma_{stop}$ is defined via the following formula:
$\sigma_{stop}=(\sigma_{max}^{\frac{1}{p_{2}}}+\frac{stop}{M}(\sigma_{min}^{\frac{1}{p_{2}}}-\sigma_{max}^{\frac{1}{p_{2}}}))^{p_{2}},$ (24)
which could be nonzero to save the one-step inference budget. By applying this special time schedule, the problems of using Free-U while preventing an increase in moving distance could be addressed, since we could only apply Free-U at timesteps that move a short distance. By leveraging the robustness of the $\beta$-VAE, the total moving distance has also been constrained.
The visual difference between our discrete method and the method proposed by the Karras’ paper is in fig.˜7
Meanwhile, since the total moving distance should be as limited as possible, we also apply the analytical first step (AFS) proposed by Dockhorn et al. (2022) during inference.
Our algorithm can be described by the algorithm˜1
@@
Data: number of inference step $N$, initiated noise $\epsilon$, stop step in Karras schedule $t_{stop}$, the total step $M$ of the original discrete method, discretion hyper parameter $p_{1}$ and $p_{2}$, U-Net augment step $t_{aug}$, ODE solver $F_{\theta}$, augment U-Net method $freeU()$, a bool value afsafs to determine whether using afs, and an analytical first step solver without using neural network $F$
@@
2
[…]
4if afsafs then
5 $ts.insert((\frac{t_{0}+t_{1}}{2}),at:1)$
@@
7
[…]
9while $t$ in tsts do
10 if $t$ equal to $t_{aug}$ then $freeU(F_{\theta})$;
11   ;
12 if afsafs and $t$ equal to 999999 then $x=F(x)$ ;
13 else $x=F_{\theta}(x)$ ;
14
@@ 3.2 Information Theory Analysis of the Different Algorithms
The experimental results in the section˜4 show that the diversity of the generative images sampled from a distillation model is the main bottleneck of its FID performance. Given that all images are generated from the same prompts, and the distillation algorithm changes the score function from pointing to the centroid of all possible images to a special image, we assume the distillation model limits the diversity of images’ low-frequency components, which is caused by modification of the score function at the start stage of the inference. Explicitly, some images with totally different semantics could have a similar, fuzzy structure with the same color block in a similar position. And the limitation on the possible fuzzy structure of images may also cause diversity problems. And we use information theory to explain why applying Free-U only in the final stage of the inference helps address this problem.
To introduce the relationship between information theory and machine learning, we should first review the historical confusion about machine learning’s generalization capability.
@@
A common point of confusion in deep learning at an early stage is how neural networks with so many parameters can still generalize (Breiman (2018)). The standard statistical model uses a limited number of parameters to concisely describe the phenomenon and capture the essence of the observed data, whereas deep learning is fundamentally different. For this problem, the study theory (Harman & Kulkarni (2012)) claims that limiting the number of functions that a model can express helps ensure that the model approaches the best functions it can have, and that it also bounds the distance between the final function a neural network study about and the average of functions.
Generally, a possible function expressed by a model is a hypothesis $\mathcal{H}$, and if the number of input data points is $N$, the number of possible hypotheses is less than $2^{N}$. According to information theory, for a dataset $\mathcal{S}$ with data $D$, the typical set that a model can access is $2^{H(D)}$ (Yeung (2012)). For example, in a high-dimensional standard Gaussian distribution, a pure black image, although with the highest possible density, is not in the typical set. Considering the amount of data in the typical set, and the mutual information between the output and the middle feature of the neural network $I(y,z)$, we can further bound the models’ capacity for potential hypotheses.
@@
Explicitly, we can consider the probability of middle feature $z$ given output $y$, and the $p(z|y)$ provides a match between $y$ and $z$. The information bottleneck theory proposed by Tishby et al. (2000) shows that $2^{H(Y|Z)}$ denotes the volume of $Y$ projected to the $Z$. Hence, $2^{I(Y,Z)}=\frac{2^{H(Y)}}{2^{H(Y|Z)}}$ further constraints the possible hypotheses by limiting the possible middle feature (Tishby et al. (2000)). As the mutual information bounds the distance between the study hypothesis and the average hypothesis, generative adversarial networks (GANs) (Goodfellow et al. (2020)), which face mode collapse (Creswell et al. (2018)) that only generate a subset of the dataset, can benefit from it by adding a mutual information constraint Jeon et al. (2021). One study (Li & Zhang (2025)) shows that this phenomenon also appears in the diffusion distillation, which consistently changes the score function and triggers an increase in mutual information. The basic idea of the proof in that study is related to $\beta$-variation auto-encoder ($\beta$-VAE). When $\beta$-VAE decreases $\beta$ to let VAE generate high-quality images, the upper bound of the mutual information constraint will be abated. Meanwhile, the peak signal-to-noise ratio (PSNR) of the encoding feature $Z$ will increase. From this perspective, the diffusion model can be deemed as a hierarchy VAE (Luo (2022)), while progressive distillation proposed by Salimans & Ho (2022) increases the mutual information by letting the model directly use $X_{\theta}({t+2})$ to replace the original output $X_{\theta}(t+1)$. Where $X_{\theta}$ is a noisy feature output by solving ODE. Hence, the PSNR of the middle output feature also increases.
@@
The diffusion ODE can be regarded as a Markov process $Z_{\theta}(T)-Z_{\theta}(T-1)-...-Z_{\theta}(t)-...Z_{\theta}(0)$, in which $Z_{\theta}(t)$ is the noisy input of the diffusion model, and we have the following eq.˜25:
$I(Z_{\theta}(0),Z_{\theta}(t))=H(Z_{\theta}(0))-H(Z_{\theta}(0)|Z_{\theta}(t))$
$\leq H(X(0))-E_{Z_{\theta}(t)}[p(Z_{\theta}(0)|Z_{\theta}(t))$
$*\log p_{\theta}(Z_{\theta}(0)|Z_{\theta}(t)))],$ (25)
in which $H(X(0))$ is the entropy of the real $X(0)$ sample from the dataset.
In $\beta$-VAE, when given the random variable $Z$, the output $Y$ is assumed to be sampled from a Gaussian distribution using $Y$ as its mean, and abating the $\beta$ makes the middle feature $Z$ more similar to $Y$, enhancing the log likelihood of the $Y$. And in the distillation algorithm, eq.˜25 shows that if we modify the $Z_{\theta}(t)$ to make it similar to the final output to enhance the log-likelihood by rectifying the trajectory, then the upper bound of the mutual information will also increase. Most distillation methods that do not increase the PSNR of the output will still face this problem. We evaluate the Precision and Recall of Distribution (PRD) (Sajjadi et al. (2018)) of flash diffusion and our sampler to further validate this claim.